Team:Groningen/Labwork/25 July 2013



Started to check the PCR reactions made on July 24th. The PCR did not went very smooth. The following combinations did show low presence of a band around 1000 bp: silk without strep + MotB and silk without strep + LytB. Although the concentration is very low, even after concentrating the samples.

Started to do the PCR gel purification kit on the Pdes and CheY. Gel extraction failed for Pdes, the end concentration of CheY is 2.9 ng/ul. Both are not useful in the rest of the project. So a new PCR is made to obtain these two PCR products. Again gel extraction is done and this time the concentration is high enough to continue. A PCR reaction is made for des UP, a gel extraction is done and the final concentration is 12 ng/ul. A PCR reaction for tet showed multiple bands, so this one needs to be purified from gel. A PCR reaction to obtain des DOWN succeeded and so did the gel extraction. The concentration is high enough to continue with. PCR reactions for CheY up, spec and CheY down unfortunately failed. These PCR reactions are made again and will be evaluated July 26th.

Did a restriction digestion with BamHI and EcoRI on the promoter in backbone BBa_k823823 obtained on July 17th, this revealed more bands on gel than expected. So it is decided to not continue with these samples.

Did a restriction digestion with EcoRI and PstI on eYFP and RFP which were obtained July 18th. The restriction digestion failed. Therefore a new restriction digestion is made. This time RFP and eYFP are run on gel as well. This revealed bands at 6000 bp for the plasmid of RFP and eYFP. Which is not correct, because bands around 700 bp are expected. Again the restriction digest failed.


Measured all the concentrations of the PCR products made by Mirjam.

Sander and Claudio

Fresh stock of IPTG 0.1M and Tryptophan 1% are prepared.
Bacillus Subtilis 168 is streaked out on LB agar.